OpenStudy (anonymous):

What are the solutions of 3x2 = 18x? A.) x = 0, x = 6 B.) x = 0, x = −6 C.) x = 3, x = 6 D.) x = 3, x = −6

OpenStudy (cggurumanjunath):

18x/(3x^2)

OpenStudy (nurali):

3x2 = 18x 3x2 -18x=0 3x(x-6)=0 3x=0 or x-6=0 x=0 or x=6

OpenStudy (anonymous):

thank u

OpenStudy (nurali):

Anytime.